how do solve the Highest Common Factor for 540,504 using Euclid method
1
Expert's answer
2012-11-22T07:42:16-0500
Decomposition into prime numbers
540=2*2*3*3*3*5
504=2*2*2*2*3*3*7
As you can see, common part in this decomposition is
2*2*3*3 = 4*9 = 36.
Hecne, HighestCommon Factor for 540,504 is 36.